Circular No. 6326

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

GRO J1744–28 T. Augusteijn and G. Van de Steene, European Southern Observatory

(ESO); F. van der Hooft, University of Amsterdam; J. van Paradijs, University of Amsterdam and University of Alabama at Huntsville; C. Kouveliotou, Universities Space Research Association and Marshall Space Flight Center (MSFC), NASA; G. J. Fishman, MSFC; and W. H. G. Lewin and R. Rutledge, Massachusetts Institute of Technology, report: “Using the ESO New Technology Telescope (NTT) and Danish 1.5-m telescope (D1.5), we have made V, R, and I images of the region around the Żytkow-Irwin star (IAUC 6314), for which we find R = 19.3 ± 0.1 (NTT, Feb. 9.323 UT), R = 19.2 ± 0.1 (D1.5, Feb. 18.362), I = 17.7 ± 0.1 (D1.5, Feb. 18.369). In a 2-min V image taken at the NTT on Jan. 21.360, we see the source just above the background, at V ∼ 21.5. In the images taken with the D1.5, we notice the presence of a second star, fainter by ∼ 2.5 mag in both R and I, offset by ∼ 2′′.1 to the west and ∼ 1′′.2 to the north. We have obtained spectra (range 550–1000 nm, resolution 2.9 nm) of the Żytkow-Irwin star, using the NTT on Feb. 9.332 and 9.347 (exposure times 15 min each). We do not find any strong spectral features in the average spectrum, with an upper limit of ∼ 0.4 nm to an emission line at Hα. From a visual comparison between the Gunn z image and the ESO K′ image (IAUC 6321), we find a good match where the Żytkow-Irwin star coincides with ‘star C’. The corresponding optical and K-band magnitudes are consistent with a source having intrinsic colors similar to those of a low-mass x-ray binary reddened by E(B − V) ∼ 3; with V ∼ 21.5 and an assumed distance of 7 kpc, the corresponding absolute visual magnitude M_{V} ∼ −2, in agreement with the value expected for its x-ray luminosity and orbital period (van Paradijs and McClintock, 1994, A.Ap. 290, 133). However, since the Żytkow-Irwin star was about equally bright in R several years ago, and transient low-mass xray binaries in outburst are expected to brighten by several magnitudes in R, it still seems possible that this source is not the true counterpart of the x-ray source. If the variable radio source (IAUC 6307) coincides with the x-ray source, we suspect that the interstellar reddening of GRO J1744–28 may substantially exceed the above value, and that its optical counterpart remains to be detected.”

1996 February 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green